A rectangles's length is 5ft more than its width. If the width is not changed and the length is increased by 3 ft, the new area will be 128ft^2. What are the dimensions of the new rectangle?
I am not sure if its this (x+5)(x+3)=128ft^2

Answer by richwmiller(17219)

x*(x+5+3)=128
x^2+8x-128=0
(x-8) (x+16) = 0
x=8
x+5=13
x+5+3=16